Darrell has 2$ less than five times as much money as Bob. Mary has as much money as Darrell and Bob put together. If all three of them together have 80$, how many money does Mary have?
Answer by stanbon(57967)

Let D,B, and M stand for the amount of money each has.
Then:
D+2=5B
M=D+B
B+D+M=80
Rearranging:
5B-D =2
B+D-M=0
B+D+M=80
Solve using Cramer or row reduction.
I get Bob has $7, Darrell has $33, and Mary has $40
Cheers,
Stan H.